# Quantum error correction for long chains of trapped ions

### v1

@@ Abstract
We propose a model for quantum computing with long chains of trapped ions and we design quantum error correction schemes for this model. The main components of a quantum error correction scheme are the quantum code and a quantum circuit called the syndrome extraction circuit, which is executed to perform error correction with this code. In this work, we design syndrome extraction circuits tailored to our ion chain model, a syndrome extraction tuning protocol to optimize these circuits, and we construct new quantum codes that outperform the state-of-the-art for chains of about $50$ qubits. To establish a baseline under the ion chain model, we simulate the performance of surface codes and bivariate bicycle (BB) codes equipped with our optimized syndrome extraction circuits. Then, we propose a new variant of BB codes defined by weight-five measurements, that we refer to as BB5 codes, and we identify BB5 codes that achieve a better minimum distance than any BB codes with the same number of logical qubits and data qubits, such as $[[30,4,5]]$ and $[[48,4,7]]$ BB5 codes. For a physical error rate of $10^{-3}$, the $[[48,4,7]]$ BB5 code achieves a logical error rate per logical qubit of $5\cdot 10^{-5}$, which is four times smaller than the best BB code in our baseline family. It also achieves the same logical error rate per logical qubit as the distance-7 surface code but using four times fewer physical qubits per logical qubit.
@@ I Introduction
To establish a baseline for the ion chain model, we consider surface codes, which achieve state-of-the-art performance for superconducting architecture [1, 2], and bivariate bicycle (BB) codes, which achieve state-of-art-performance when long-range gates are available [47]. We simulate the performance of these codes with a syndrome extraction circuit optimized for the ion chaim model. Our baseline family includes surface codes with distance $3,5$ and $7$. The number of BB codes with up to 50 data qubits is massive and some of them have poor parameters. Therefore, we do not simulate all of them. Instead, we simulate a BB code with optimal minimum distance $d$ for each achievable pair of $[[n,k]]$ within the BB code family for $n\leq 50$.
Finally, we propose a new variant of BB codes defined by weight-5 stabilizer generators, in contrast to the weight-6 stabilizer generators of the original BB codes. We refer to these new codes as BB5 codes. To avoid any confusion, in the rest of this paper we use the term BB6 codes for the original BB codes of [47]. We identify two BB5 codes with parameters $[[30,4,5]]$ and $[[48,4,7]]$ that achieve larger code distance than any BB6 code with the same parameters $[[n,k]]$. For comparison BB6 codes only achieve $[[30,4,4]]$ and $[[48,4,6]]$. Our simulations show that, under the ion chain model, the $[[48,4,7]]$ BB5 code achieves a logical error rate per logical qubit $4$ times smaller than the best BB6 codes with code length up to $50$. Moreover, it reaches the same logical error rate per logical qubit as the distance-7 surface code while using $4$ times fewer physical qubits per logical qubit.
@@ II The ion chain model
Finally, we assume that all the operations have the same duration, except measurements which take $\tau_{m}$ times longer. This long measurement time translates to more idle noise during a measurement. We consider a regime with $\tau_{m}p<<100$ and we set the idling noise rate during a measurement to $\tau_{m}p/100$. In our simulations, we use $\tau_{m}\approx 30$.
@@ IV Improved codes for the ion chain model
Rows of ${\bf{H}}_{X}$ and ${\bf{H}}_{Z}$ are respectively the indicator vectors of the $X$ and $Z$ stabilizer generators of the codes [51, 52]. Each $A_{i}$ is an $(m\ell)\times(m\ell)$ permutation matrix of the form $Q_{\ell}^{u}\otimes Q_{m}^{v}$ with $u\in\{0,1,\dots,\ell-1\}$ and $v\in\{0,1,\dots,m-1\}$. Each BB5 code is specified by the values of $\ell,m$ and the matrices $A_{1},\dots,A_{5}$. Table 2 presents two instances of BB5 codes, found by exhaustive search, achieving a larger minimum distance than any BB6 codes with the same parameter $[[n,k]]$.
@@
The logical error rate is estimated with a Monte Carlo simulation using Stim [53], the BP-OSD decoder [54, 55, 56] for BB5 and BB6 codes, and PyMatching 2 [57] for surface codes. For all the simulations in this paper, the relative error bar is smaller than $6\%$.

### v2

@@ Abstract
We propose a model for quantum computing with long chains of trapped ions and we design quantum error correction schemes for this model. The main components of a quantum error correction scheme are the quantum code and a quantum circuit called the syndrome extraction circuit, which is executed to perform error correction with this code. In this work, we design syndrome extraction circuits tailored to our ion chain model, a syndrome extraction tuning protocol to optimize these circuits, and we construct new quantum codes that outperform the state-of-the-art for chains of about $50$ qubits. To establish a baseline under the ion chain model, we simulate the performance of surface codes and bivariate bicycle (BB) codes equipped with our optimized syndrome extraction circuits. Then, we propose a new variant of BB codes defined by weight-five measurements, that we refer to as BB5 codes and we identify BB5 codes that achieve a better minimum distance than any BB codes with the same number of logical qubits and data qubits, such as a $[[48,4,7]]$ BB5 code. For a physical error rate of $10^{-3}$, the $[[48,4,7]]$ BB5 code achieves a logical error rate per logical qubit of $5\cdot 10^{-5}$, which is four times smaller than the best BB code in our baseline family. It also achieves the same logical error rate per logical qubit as the distance-7 surface code but using four times fewer physical qubits per logical qubit.
@@ I Introduction
To establish a baseline for the ion chain model, we consider surface codes, which achieve state-of-the-art performance for superconducting architecture [1, 2], and bivariate bicycle (BB) codes, which achieve state-of-art-performance when long-range gates are available [47]. We simulate the performance of these codes with a syndrome extraction circuit optimized for the ion chain model. Our baseline family includes surface codes with distance $3,5$ and $7$. The number of BB codes with up to 50 data qubits is massive and some of them have poor parameters. Therefore, we do not simulate all of them. Instead, we simulate a BB code with optimal minimum distance $d$ for each achievable pair of $[[n,k]]$ within the BB code family for $n\leq 50$.
Finally, we study a variant of BB codes defined by weight-5 stabilizer generators, in contrast to the weight-6 stabilizer generators of the original BB codes. We refer to these new codes as BB5 codes. To avoid any confusion, in the rest of this paper we use the term BB6 codes for the original BB codes of [47]. We identify two BB5 codes with parameters $[[30,4,5]]$ and $[[48,4,7]]$ that achieve larger code distance than any BB6 code with the same parameters $[[n,k]]$. For comparison BB6 codes only achieve $[[30,4,4]]$ and $[[48,4,6]]$. The BB5 codes we consider belong to the broader families of tensor-product generalized bicycle codes [50] and multivariate bicycle codes proposed in [51]. In [51] the authors previously discovered a code with the same parameters as our $[[30,4,5]]$ BB5 code by searching over a set of codes that is a subset of BB5 codes.
Our simulations show that, under the ion chain model, the $[[48,4,7]]$ BB5 code achieves a logical error rate per logical qubit $4$ times smaller than the best BB6 codes with code length up to $50$. Moreover, it reaches the same logical error rate per logical qubit as the distance-7 surface code while using $4$ times fewer physical qubits per logical qubit.
@@ II The ion chain model
Finally, we assume that all the operations have the same duration, except measurements which take $\tau_{m}$ times longer. This long measurement time translates to more idle noise during a measurement. We consider a regime with $\tau_{m}p<<100$ and we set the idling noise rate during a measurement to $\tau_{m}p/100$. In our simulations, we use $\tau_{m}=30$.
@@ IV Improved codes for the ion chain model
Rows of ${\bf{H}}_{X}$ and ${\bf{H}}_{Z}$ are respectively the indicator vectors of the $X$ and $Z$ stabilizer generators of the codes [53, 54]. Each $A_{i}$ is an $(m\ell)\times(m\ell)$ permutation matrix of the form $Q_{\ell}^{u}\otimes Q_{m}^{v}$ with $u\in\{0,1,\dots,\ell-1\}$ and $v\in\{0,1,\dots,m-1\}$. Each BB5 code is specified by the values of $\ell,m$ and the matrices $A_{1},\dots,A_{5}$. Table 2 presents two instances of BB5 codes, found by exhaustive search, achieving a larger minimum distance than any BB6 codes with the same parameter $[[n,k]]$. The construction of ${\bf{H}}_{X}$ and ${\bf{H}}_{Z}$ from the matrices $A_{i}$ is identical to Prop. 1 in [51], however we allow for matrices $A_{i}=Q_{\ell}^{u}\otimes Q_{m}^{v}$ whereas the previous construction is restricted to $u=v$. Note the construction of our $[[48,4,7]]$ code uses $u\neq v$ (see Table 2).
@@
The logical error rate is estimated with a Monte Carlo simulation using Stim [55], the BP-OSD decoder [56, 57, 58] for BB5 and BB6 codes, and PyMatching 2 [59] for surface codes. We use BP-OSD with parameters max_bp_iters = 10_000, bp_method = "min_sum", osd_order = 5, osd_method = "osd_cs" . For all the simulations in this paper, the relative error bar is smaller than $6\%$.